OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \frac{ \sqrt[6]{x}-1 }{ \sqrt[3]{x}-1 }\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \frac{ \sqrt[6]{x}-1 }{ \sqrt[3]{x}-1 }\]

OpenStudy (anonymous):

please help